OpenStudy (anonymous):

6x-3x+10=-2 step by step plz

OpenStudy (agreene):

6x-3x+10=-2 combine like terms (6-3)x+10=-2 3x+10=-2 subtract 10 from both sides 3x=-12 divide by 3 x=-12/3 evaluate fraction x=-4